# HealthSLM-Bench: Benchmarking Small Language Models for Mobile and Wearable Healthcare Monitoring

### v2

@@ 5 Results and Discussion
x PMData GLOBEM AW-FB Model STRS ($\downarrow$) READ ($\downarrow$) FATG ($\uparrow$) SQ ($\downarrow$) ANX ($\downarrow$) DEP ($\downarrow$) ACT ($\uparrow$) CAL ($\downarrow$) LLMs (ZS) MedAlpaca 0.76 2.18 46.8 0.68 1.23 0.89 21.7 35.0 PMC-Llama 1.33 4.83 0.00 2.25 2.33 2.23 43.4 Asclepius 0.43 1.44 27.3 0.45 0.82 1.10 28.9 ClinicalCamel 0.40 2.11 58.1 0.37 0.97 0.79 16.3 43.4 Flan-T5 0.36 1.82 56.8 0.56 2.84 2.89 23.4 66.0 Palmyra-Med 0.83 5.01 43.5 0.44 2.07 1.99 29.7 75.3 Llama 2 0.57 2.86 41.2 0.89 1.19 1.23 BioMedGPT 0.37 2.12 61.2 0.41 0.95 0.85 12.2 BioMistral 0.55 2.12 56.6 0.45 0.90 18.4 41.0 GPT-3.5 2.38 70.8 0.87 13.8 36.4 GPT-4 2.22 72.2 0.73 22.6 75.2 Gemini-Pro 0.79 1.69 34.0 0.78 1.03 0.95 17.7 31.4 Mean 0.64 2.56 41.54 0.60 1.43 1.44 19.53 47.60 SLMs (ZS) Gemma-2-2b-it 0.72 2.07 52.84 0.47 0.91 0.53 - 105.12 Phi-3-mini-4k 0.46 1.52 62.88 0.48 1.08 1.26 17.39 93.80 SmolLM-1.7B 1.42 2.99 11.04 1.00 2.59 2.87 21.74 277.21 Qwen2-1.5B 0.40 2.03 63.21 0.45 1.42 1.65 14.05 185.22 TinyLlama-1.1B 0.43 2.06 51.17 0.47 2.40 2.58 19.73 198.72 Llama-3.2-1B 0.40 1.87 63.79 0.69 1.51 1.85 11.71 280.32 Llama-3.2-3B 0.67 2.24 40.80 0.47 1.26 0.75 15.72 19.70 Phi-3.5-mini 0.41 2.34 61.20 0.46 0.88 0.84 15.38 56.75 Qwen2.5-1.5B 0.56 2.25 62.88 0.93 1.37 1.63 15.72 72.20 Mean 0.61 2.15 52.20 0.60 1.49 1.55 16.40 143.23
@@ 5.1 Overall Performance
As shown in Table 5, both SLMs and SOTA LLMs [23] are instruction-tuned, yet SLMs outperform LLMs in tasks such as fatigue and calorie estimation. Specifically, SLMs achieve much lower mean values for fatigue (46.1 for SLMs vs. 52.4 for LLMs) and for calorie estimation error (7.57 for SLMs compared to 41.6 for LLMs), demonstrating their superior accuracy for these important health measures. Although LLMs perform slightly better in stress, readiness, and activity prediction—with lower mean values for stress (0.44 for LLMs vs. 0.57 for SLMs), readiness (1.48 vs. 1.68), and higher mean values for activity (28.2 vs. 21.8)—these differences are relatively modest compared to the clear advantages of SLMs in fatigue and calorie estimation. For other tasks such as sleep quality, anxiety, and depression, both SLMs and LLMs show similar performance, with only minor differences in mean values. Notably, SLM models like TinyLlama-1.1B and Phi-3-mini-4k stand out for their strong and consistent results across multiple tasks. For the less-performing cases (e.g., activity, anxiety and sleep quality) of SLMs, we observed that SLMs tend to predict only the majority classes without attempting to predict other, weaker classes (i.e., class-imbalance bias; cf. Appendix), causing the model to become stuck at sub-optimal performance on thoes tasks. However, we note that the performance differences on these tasks, compared with LLMs, are acceptable.

### v3

@@ 5 Results and Discussion
PMData GLOBEM AW-FB
LLMs (ZS) MedAlpaca 0.76 2.18 46.8 0.68 1.23 0.89 21.7 35.0
PMC-Llama 1.33 4.83 0.00 2.25 2.33 2.23 43.4
Asclepius 0.43 1.44 27.3 0.45 0.82 1.10 28.9
ClinicalCamel 0.40 2.11 58.1 0.37 0.97 0.79 16.3 43.4
Flan-T5 0.36 1.82 56.8 0.56 2.84 2.89 23.4 66.0
Palmyra-Med 0.83 5.01 43.5 0.44 2.07 1.99 29.7 75.3
Llama 2 0.57 2.86 41.2 0.89 1.19 1.23
BioMedGPT 0.37 2.12 61.2 0.41 0.95 0.85 12.2
BioMistral 0.55 2.12 56.6 0.45 0.90 18.4 41.0
GPT-3.5 2.38 70.8 0.87 13.8 36.4
GPT-4 2.22 72.2 0.73 22.6 75.2
Gemini-Pro 0.79 1.69 34.0 0.78 1.03 0.95 17.7 31.4
Mean 0.64 2.56 41.54 0.60 1.43 1.44 19.53 47.60
SLMs (ZS) Gemma-2-2b-it 0.72 2.07 52.84 0.47 0.91 0.53 - 105.12
Phi-3-mini-4k 0.46 1.52 62.88 0.48 1.08 1.26 17.39 93.80
SmolLM-1.7B 1.42 2.99 11.04 1.00 2.59 2.87 21.74 277.21
Qwen2-1.5B 0.40 2.03 63.21 0.45 1.42 1.65 14.05 185.22
TinyLlama-1.1B 0.43 2.06 51.17 0.47 2.40 2.58 19.73 198.72
Llama-3.2-1B 0.40 1.87 63.79 0.69 1.51 1.85 11.71 280.32
Llama-3.2-3B 0.67 2.24 40.80 0.47 1.26 0.75 15.72 19.70
Phi-3.5-mini 0.41 2.34 61.20 0.46 0.88 0.84 15.38 56.75
Qwen2.5-1.5B 0.56 2.25 62.88 0.93 1.37 1.63 15.72 72.20
Mean 0.61 2.15 52.20 0.60 1.49 1.55 16.40 143.23
@@ 5.1 Overall Performance
As shown in Table 5, both SLMs and SOTA LLMs [23] are instruction-tuned, yet SLMs outperform LLMs in tasks such as fatigue and calorie estimation. Specifically, SLMs achieve much lower mean values for fatigue (46.1 for SLMs vs. 52.4 for LLMs) and for calorie estimation error (7.57 for SLMs compared to 41.6 for LLMs), demonstrating their superior accuracy for these important health measures. Although LLMs perform slightly better in stress, readiness, and activity prediction—with lower mean values for stress (0.44 for LLMs vs. 0.57 for SLMs), readiness (1.48 vs. 1.68), and higher mean values for activity (28.2 vs. 21.8)—these differences are relatively modest compared to the clear advantages of SLMs in fatigue and calorie estimation. For other tasks such as sleep quality, anxiety, and depression, both SLMs and LLMs show similar performance, with only minor differences in mean values. Notably, SLM models like TinyLlama-1.1B and Phi-3-mini-4k stand out for their strong and consistent results across multiple tasks. For the less-performing cases (e.g., activity, anxiety and sleep quality) of SLMs, we observed that SLMs tend to predict only the majority classes without attempting to predict other, weaker classes (i.e., class-imbalance bias; cf. Appendix), causing the model to become stuck at sub-optimal performance on those tasks. However, we note that the performance differences on these tasks, compared with LLMs, remain relatively minor.